OpenStudy (anonymous):

The school purchased six dozen new volleyballs.it distributed all all but eight balls. somehow the 7th and 8th grade boys gt the same numbers of balls-one more then the 7th grade girls.the 8thgrade girls got 5 more balls either boys team received. how many volley ball did each team receive?

OpenStudy (anonymous):

your question seems to be missing some words... It looks sort of like: 7th grade boys = 8th grade boys = x = 7th grade girls + 1 8th grade girls = 5 + 7th grade boys = 5 + 8th grade boys Question: solve for 7th grade boys, 8th grade boys, 7th grade girls, 8th grade girls